OpenStudy (anonymous):

what is the value of the following expression? tan(-5pi/6)

OpenStudy (anonymous):

\[\tan \left(-\frac{ 5\pi }{ 6 } \right) \]

OpenStudy (anonymous):

Starting from zero on the unit circle, rotate clockwise (5/6)π You will end up with a reference angle of π/6 in quadrant III where tan>0 tan(-5π/6)=tan(π/6)=√3/3